OpenStudy (anonymous):

What is the equation of the line that passes through (4, 2) and is parallel to 3x – 2y = –6

OpenStudy (texaschic101):

first we will turn the equation 3x - 2y = -6 into y = mx + b form...can you do that ?

OpenStudy (anonymous):

yea

OpenStudy (texaschic101):

show me what you get when you do that

OpenStudy (texaschic101):

or do you need help with that ?

OpenStudy (anonymous):

im confused

OpenStudy (anonymous):

do (4,2) go with it?

OpenStudy (texaschic101):

3x - 2y = -6 -- subtract 3x from both sides -2y = -3x - 6 -- divide both sides by -2 y = 3/2x + 3 do you understand this so far ?...we will get to (4,2) in a minute

OpenStudy (texaschic101):

what we are doing is turning the equation into y = mx + b form to find the slope, which is in the m position

OpenStudy (anonymous):

ok i get it now

OpenStudy (texaschic101):

y = mx + b y = 3/2x + 3 so our slope in this equation is 3/2 understand so far ?

OpenStudy (texaschic101):

if you have questions, ask them...if not I will continue...let me know

OpenStudy (anonymous):

im listening

OpenStudy (texaschic101):

The reason we need to find the slope is because a parallel line will have the same slope. So we know that the slope is 3/2, but we do not know the y intercept, which is in the b position of y = mx + b. Now we take y = mx + b slope(m) = 3/2 here is where your points come in...(4,2)...x = 4 and y = 2 now we sub...we are looking for b, the y intercept 2 = 3/2(4) + b 2 = 6 + b -- subtract 6 from both sides 2 - 6 = b -4 = b so your parallel equation is : y = 3/2x - 4 understand ?

OpenStudy (texaschic101):

now thats slope intercept form...or do you need it in standard form Ax + By + C ?

OpenStudy (texaschic101):

oops..typo..Ax + By = C

OpenStudy (anonymous):

i have it thank you can you help me with another one?

OpenStudy (texaschic101):

I think I can...we will see...what ya got ?

OpenStudy (anonymous):

The graph of which of the following will be parallel to the graph of 4x – 3y = –12?

OpenStudy (anonymous):

4x-3y=-12

OpenStudy (texaschic101):

parallel line will have the same slope, but different y intercepts remember, in y = mx + b form, the slope is in the m position and the y intercept is in the b position. 4x - 3y = -12 -3y = -4x - 12 y = 4/3x + 4 .....slope is 4/3 and y intercept is -12 4x - 3y = -12 -3y = -4x - 12 y = 4/3x + 4.....this is not parallel....the slopes and the y intercepts are the same, meaning it is the same line. do you have more answer choices ?

OpenStudy (anonymous):

B. 6x – 4y = –8 C. D. 4x – 2y = –12

OpenStudy (anonymous):

A. y=4/3x -3/2 c. y=3/4x+1

OpenStudy (texaschic101):

ok...in the original problem we found the slope to be 4/3 and the y intercept 4. we need the same slope, but different y intercepts for the lines to be parallel. Now look at this.. y = 4/3x - 3/2....same slope, but different y intercepts...thats what we want and this is your answer

OpenStudy (anonymous):

Thank you

OpenStudy (texaschic101):

sure thing :)